OpenStudy (anonymous):

You place a cup of 205oF coffee on a table in a room that is 72oF, and 10 minutes later, it is 195oF. Approximately how long will it be before the coffee is 180oF? Use Newton's law of cooling: A. 1 hour B. 25 minutes C. 45 minutes D. 15 minutes

OpenStudy (anonymous):

work always with the differences between the heated coffee and the room temp first number we need is \(205-72=133\)

OpenStudy (anonymous):

second number is \(195-72=123\)

OpenStudy (anonymous):

this takes 10 minutes to occur, so you can model this as \[133\left(\frac{123}{133}\right)^{\frac{t}{10}}\]

OpenStudy (anonymous):

since \(180-72=108\) set \[108=133\left(\frac{123}{133}\right)^{\frac{t}{10}}\] and solve for (t\)